# The Face of Populism: Examining Differences in Facial Emotional Expressions of Political Leaders Using Machine Learning

### v4

@@ 4 Results
Results suggest that negative emotions are more frequently expressed in the entire sample, with anger (0.227) as the emotion having the highest average score, followed by sadness (0.206) and fear (0.127). As a consequence, the average proportion of negative emotions is greater than 0.5, which means that on average negative emotions are detected more frequently than positive emotions (e.g., happiness) and neutral expressions combined. There are clear issues with the detection of disgust and surprise, where a low maximum level suggests low positive predictive value for these emotions. However, our focus is on differences in the expression of negative emotions and in scores of neutral states. Aggregating negative emotions enables us to surpass the false positive issues within the group of negative emotions (e.g., expression of fear being classified as sadness, disgust as anger, Khaireddin & Chen 2021) and minimizes the impact of the low positive predictive value of disgust and surprise.
@@
The mean average score of negative emotions in the populist group is 0.616 (95% CI 0.584–0.649), compared to the 0.500 mean average score for pluralists (95% CI 0.464–0.537). The difference between the two groups is statistically significant ($t=4.691$, $p=0.000005$, $d=0.66$ moderate effect size). This finding suggests that pluralists generally tend to have an average score of negative emotions around 0.5, which roughly translates to the situation where negative emotion is the prevailing expression in one half of the video (in fact, that percentage is 46.1% compared to 66.7% for populists; see Table 3 in the Supplement). On the other hand, for populists, more than one half of each video where they appear is dominated by negative emotions.
@@ 5 Discussion
The present study examines facial emotional expressions of populist and non-populist political leaders across fifteen countries using a deep-learning-based computer-vision algorithm to investigate differences in their expressions of negative emotions (anger, disgust, fear, and sadness) and neutral stances, i.e., facial expressions which lack emotional disposition. Our findings indicate that populist leaders tend to express negative emotions more frequently than non-populist leaders, and conversely, that neutral expressions are less common in their public performances than they are in those of more conventional, pluralist political figures. The significance and the effect size of these differences are higher compared to differences between ideological groups (see Supplement E). These findings align with the theoretical framework on the characteristics of affective populist communication (Bonansinga 2020b, Salmela & von Scheve 2018, Engesser et al. 2017), as well as with previous, mostly text-based research which indicates that populist leaders employ language which is associated with negative emotion at a higher rate than mainstream, pluralist politicians (Breeze 2019, Caiani & Di Cocco 2023, Macagno 2022, Martella & Bracciale 2022, Widmann 2021).
@@
Overall, the results provide new evidence of a style of affective, non-verbal communication more common to political leaders whose rhetoric is populist-leaning, and thus further substantiate the idea that populism is performative in nature Bucy et al. (2020), Moffitt (2022b). As such, because it is subversive in relation to established political conventions, the populist performance is especially well-suited for the environment of social media where the principle of “any press is good press” often applies—that is, the quantity of audience attention bears more weight than its quality (Wells et al. 2020). On one hand, the populist use of negative emotional displays in video may therefore be a mechanism through which they maintain relevance, considering that controversial content online ordinarily generates more engagement among consumers (Larsson 2022). On the other, populists’ facial expressions can provide affective heuristics for audiences, influencing their evaluation of candidates’ political competence (Boussalis & Coan 2021) and allowing them to quickly reduce complex political issues into simple, emotional interpretations by visually framing them as negative. Further analysis of the role of non-verbal cues in political communication—such as facial expressions of emotion—would be beneficial to gaining a better understanding of the populist performance as a whole.
@@
Firstly, the sample of YouTube videos is not globally representative and does not provide ample coverage of the global political spectrum in regard to the use of populist or pluralistic rhetoric. As we have shown in the supplementary analysis (see Supplement C: Country-level differences), the differences between populist and pluralist leaders are not consistent in all investigated countries. Larger research infrastructure is required to process a more substantial number of videos, leaders and countries and the current computational toolset needs to be extended to enable seamless automatic processing of videos. This also relates back to the gender imbalance of the current sample. While it is reasonable to hypothesize that gender influences emotional expression among political leaders (Boussalis et al. 2021), in thiscase, including more women without significantly increasing the overall size of the sample and the number of leaders would mean focusing only on countries where there is at least one female leader of a non-fringe political party. This creates the risk of introducing bias towards Western countries in the sample and disregarding the global scope.
@@
Secondly, a key component of our analysis is the measure of populist rhetoric supplied by the Global Party Survey database (Norris 2020) which significantly limits the timeframe of the analysis (no leaders, parties and videos after 2019 were taken into consideration), but also imposes the inherent limits of an expert survey. While the GPS is the only available source of multi-dimensional measurements of populist rhetoric on a global scale, the experts surveyed may operate within different understandings of what populist rhetoric is. This is particularly challenging in developing countries, where fewer experts answered the survey, as well as in autocratic regimes where experts may not have been able to provide a fair assessment of the (ruling) political parties (or refused to do so). By focusing on large countries with free elections, we tried to mitigate some of the risks, but there is the possibility that some measurements used in this study reflect expert biases, especially since we used a single, aggregate measure of populist rhetoric. Future studies should aim for more granularity and robustness in the measurement of populist rhetoric, which may be obtained by following a text-as-data approach and obtaining the measures using computational and machine learning tools, where expert survey scores provide external validation (Jankowski & Huber 2023).
@@
Furthermore, since we lacked the resources necessary to process every political speech from each party’s official YouTube channel, the first criterion of selection is the view count of the videos. Our expectation is that this metric serves as a proxy for the relevance of the speech content to a wider audience over a longer period of time. However, since we have no information about the audience, it is possible we are observing a demand-side effect, where audiences who consume populist videos might show a stronger preference for those that prominently feature negative emotions. This could introduce a specific bias in the sample.
One of the major limitations of our current workflow is that the video-processing algorithm does not focus on one singular face and does not consistently label the same faces between frame jumps, which requires researchers to select videos with consistent framing and a small number of different faces. Solving this problem will allow for faster and more reliable collection of larger samples.
@@
The next major limitation is a result of the machine learning model deployed for emotion recognition. We conducted a small validation study to compare the performance of our model with the consensus of 5 human coders. Although results show evidence of moderate agreement between humans and the machine learning model, there is high inconsistency between human coders even for the simple task of assessing whether a given facial expression shows positive emotion, negative emotion, or a neutral expression. There is an urgent need for large-scale image annotation studies with expert coders if we are to assess the performance and accuracy of FER models in the future, especially on difficult “in the wild” image datasets such as ours. However, it is worth noting that in our main analysis we focus on averages across 300 frames of each video, making our analysis robust to misclassification of a smaller number of individual frames. That being said, we believe that fair agreement between the human majority vote and machine learning model does not compromise the validity of our findings.
@@
Overall confidence in the presented results would surely be higher if we based our analysis on a computer-vision approach which was shown to have higher accuracy, at least on FER2013 data. One of the possible alternatives would be to follow recent work (Bossetta & Schmøkel 2023, Schmøkel & Bossetta 2022) and use proprietary API such as Amazon’s Rekognition. Nonetheless, the aim of this paper was to introduce an open-source approach focusing on free computing resources and the deployment of machine learning algorithms on systems that are accessible to individual researchers in social sciences, including graduate students without significant institutional support for computational work.
This study was focused on exploring basic mean-level differences between the given groups. It is worth noting that the full output of our emotion detection workflow includes time series data which are suitable for multivariate analysis. Even though time series are compositional in nature, they can be transformed (Greenacre 2021) so that traditional time series methods can be applied to explore affective dynamics (Ryan et al. 2023), find additional differences in relation to other factors (such as type of public address, the proximity of elections, ideology,etc.) and cluster political leaders according to patterns of correlations in time-series. Furthermore, multivariate change point detection methods (James & Matteson 2015) could be used to identify the key moments in speech where significant changes in emotional expression occur, which can further be linked to textual analyses of transcripts. This would allow researchers to identify the rhetorical elements which trigger switches in the emotional dynamics of leaders’ political addresses.
@@ 6 Data availability
Data files are stored as CSV files. Processed data used for both the main and supplementary analyses can be found at: https://github.com/atomashevic/face-of-populism. Repository contains a CSV file with a list of URLs of all YouTube videos used in the analysis as well as Python code that can be used to reproduce the CSV files with processed data. CSV file extracted from GPS data is available in the repistory and the original GPS dataset is available at: https://dataverse.harvard.edu/dataverse/GlobalPartySurvey.
Code used for image annotation and the resulting dataset can be found at: https://github.com/atomashevic/fop-annotation.
@@ 7 Acknowledgements
The authors did not receive any financial support for this research. The annotation of images was made possible thanks to the dedicated work and effort of our undergraduate students, A. Nadj, M. Milanko, and A. Mišić, who volunteered as coders.
Appendix A Supplementary: Alternative frame processing approach
In our main analysis we presented the results based on the frame-processing approach of selecting 300 uniformly distributed frames from each video. An alternative, which yields roughly the same number of frames per video (273.03 on average), is to select and process every 50th frame of each given video. With this approach, we arrive at 209 successfully processed videos—six more than in the main analysis, because we find more consistent face selection in several videos when the algorithm receives this type of input. For shorter videos, this reduces the probability that, when selecting exactly 300 frames, a significant proportion of these frames captures cutaways to other people or the audience instead of the leader delivering their speech. Pluralist leaders are present in 101 videos (31 strongly pluralist and 70 moderately pluralist) and populist leaders are present in the remaining 108 videos (70 strongly populist and 38 moderately populist). Descriptive statistics for emotion variables obtained from this frame processing approach are presented in Table 2.
Mean SD Min. Median Max.
Anger 0.223 0.130 0.031 0.195 0.728
Disgust 0.003 0.006 0.000 0.001 0.051
Fear 0.122 0.087 0.015 0.093 0.530
Happiness 0.084 0.099 0.001 0.053 0.594
Sadness 0.204 0.109 0.021 0.183 0.573
Surprise 0.030 0.037 0.000 0.018 0.275
Neutral 0.332 0.167 0.021 0.309 0.812
Negative emotions 0.552 0.182 0.147 0.558 0.965
Table 2: Descriptive statistics for extracted emotions for the entire sample, $N=209$ videos
These results are quite similar to those presented in Table 1, with no major deviations in terms of central tendency or variability of the 8 variables. Anger, sadness, and fear are the most common emotions and negative emotions have an average score higher than 0.5.
Again, we observe statistically significant differences in the means of negative emotions and neutral expression between populists and pluralists (negative emotions: $t=3.742$, $p=0.00024$; neutral expression: $t=-3.222,p=0.00148$). The sample mean differences are smaller than in the main analysis, but with no changes in the statistical significance (negative: 0.116 in main analysis, 0.091 alternative; neutral: -0.075 in main analysis, - 0.074 in alternative). The distribution of average scores is shown in Figure 5.
Figures 6 and 7 illustrate the differences in the distribution of mean emotion scores for four subsamples according to the degree of pluralist/populist rhetoric. The results of one-way ANOVA remain in accordance with the main analysis, but with weaker effect size in the case of negative emotions. The differences between the group means are statistically significant both in the case of negative emotions ($F=4.886$, $p=0.002712$) with moderate effect size measured by eta-squared (0.073), and neutral expression ($F=5.180$ $p=0.001846$) with moderate effect size (0.077).
Figure 6: Raincloud plot showing the distribution of average scores of negative emotions across four degrees of populism. Every data point is presented as a colored dot and above them boxplots and density plots are displayed. The dashed vertical bar indicates the score of 0.5
The results from Tukey HSD post-hoc tests reveal significant differences between leaders coming from parties employing strong populist rhetoric and leaders coming from pluralist parties (mean difference 0.1298 $p=0.0057$ in comparison with strong pluralists, and mean difference 0.0961 $p=0.0166$ in comparison with moderate pluralists). For neutral expression we again find the opposite case, as strong pluralists differ from both groups of populists (mean difference 0.138 $p=0.0037$ for moderate populists, and mean difference 0.1131 $p=0.0122$ for strong populists).
Overall, we conclude that using an alternative frame processing approach does not change the outcomes of statistical tests, but the reduced effect size in the case of negative emotions and overall shrinkage of the absolute differences between groups should certainly be noted.
Appendix B Supplementary: Dominant emotions per frame
Another way of conceptualizing differences in emotional expression is to look at the proportions of frames in which a certain emotion or type of emotional expression is dominant among the groups (Table 3).
[…]
Main analysis
Pluralists 0.162 0.001 0.094 0.072 0.123 0.019 0.428 0.461
Populists 0.262 0.004 0.084 0.062 0.234 0.010 0.345 0.667
Supplementary
Pluralists 0.176 0.000 0.096 0.073 0.132 0.020 0.438 0.484
Populists 0.248 0.000 0.080 0.076 0.225 0.011 0.358 0.640
Table 3: Mean proportion of all frames in which emotion or group of emotions has the highest score, and mean proportion of frames in which negative emotions have a sum score $\geq 0.5$, per category of leader and split by frame processing approach.
Appendix C Supplement C: Country-level comparison
The sample of YouTube videos and leaders who appear in them is not representative for any country included. We have selected only a few political parties and deliberately excluded small parties from the analysis. Nonetheless, here we provide a brief overview of raw and mean scores of negative emotions and neutral expression for every country in the sample, split by populist and pluralist leaders. Both Figure 8 and 9 show that the differences between populists and pluralists are not always consistent between the countries.
Figure 8: Mean and raw scores of negative emotions for every country (ISO country codes on x-axis) in the sample and for populist and pluralist leaders. Transparent points are raw scores, thicker points are means for every group.
For example, in the case of Serbia, Croatia and South Africa, pluralists show more negative emotions (average score $>0.5$ in all three cases). On the other hand, in Serbia, Croatia, Australia, the Netherlands and Brazil, we have cases where populists are on average more neutral in their public appearances.
Figure 9: Mean and raw scores of neutral expression for every country (ISO country codes on x-axis) in the sample and for populist and pluralist leaders. Transparent points are raw scores, thicker points are means for every group.
Larger samples with more coverage of political parties and movements are needed in order to investigate whether these differences and inconsistencies are consequences of sampling variability or true effects.
Appendix D Supplementary: Political leaders
Table 4: Information about political leaders present in the YouTube video dataset, the parties they are affiliated with, party or coalition acronyms from the GPS dataset, and populism and ideology categories from the GPS dataset (TYPE_POPULISM and TYPE_VALUES variables)
Country Leader Party Acronym Populism Category Ideology Category
Argentina Alberto Fernández Justicialist Party PJ Strongly populist Right Conservative
Mauricio Macri Republican Proposal PRO-UCR (coalition) Moderately pluralist Right Conservative
Sergio Massa Renewal Front 1País (coalition) Moderately populist Right Conservative
Australia Warren Truss National Party of Australia NAT Moderately populist Right Conservative
Bill Shorten Australian Labor Party ALP Strongly pluralist Left Liberal
Scott Morrison Liberal Party of Australia LP Moderately pluralist Right Conservative
Brazil Simone Tebet Brazilian Democratic Movement MDB Moderately pluralist Right Conservative
Luiz Inácio Lula da Silva Workers’ Party PT Moderately pluralist Left Liberal
Jair Bolsonaro Social Liberal Party PSL Strongly populist Right Conservative
Croatia Andrej Plenković Croatian Democratic Union HDZ Moderately populist Left Conservative
Zoran Milanović Social Democratic Party SPH (coalition) Moderately pluralist Left Liberal
Ivan Vilibor Sinčić Key of Croatia (formerly named Human Shield) OOP (coalition) Strongly populist Right Conservative
France Marine Le Pen National Rally FN Strongly populist Right Conservative
Emmanuel Macron Renaissance (formerly named En Marche) REM Moderately pluralist Right Liberal
François Bayrou Democratic Movement MoDem Strongly pluralist Right Liberal
Hungary Viktor Orbán Fidesz Fidesz Strongly populist Right Conservative
Ferenc Gyurcsány Democratic Coalition DK Strongly pluralist Right Liberal
Péter Jakab Jobbik Jobbik Strongly populist Right Conservative
India Narendra Modi Bharatiya Janata Party BJP Strongly populist Right Conservative
Sonia Gandhi Indian National Congress INC Moderately pluralist Left Liberal
Mamata Banerjee All India Trinamool Congress AITC Moderately populist Left Conservative
Italy Matteo Salvini Lega Nord CD (coalition) Strongly populist Right Conservative
Enrico Letta Democratic Party CCS (coalition) Strongly pluralist Left Liberal
Beppe Grillo Five Star Movement M5S Strongly populist Left Conservative
Japan Shinzo Abe Liberal Democratic Party LDP Moderately populist Right Conservative
Natsuo Yamaguchi Komeito NKP Strongly pluralist Left Conservative
Yukio Edano Constitutional Democratic Party of Japan CDP Moderately pluralist Left Liberal
Netherlands Geert Wilders Party for Freedom PVV Strongly populist Left Conservative
Alexander Pechtold Democrats 66 D66 Strongly pluralist Right Liberal
Mark Rutte People’s Party for Freedom and Democracy VVD Moderately pluralist Right Liberal
Serbia Aleksandar Vučić Serbian Progressive Party SNS Strongly populist Right Conservative
Zoran Lutovac Democratic Party DS Moderately pluralist Left Liberal
Saša Radulović Enough is Enough DJB Moderately pluralist Right Liberal
South Africa Julius Malema Economic Freedom Fighters EFF Strongly populist Left Conservative
Jacob Zuma African National Congress ANC Moderately populist Left Conservative
John Steenhuisen Democratic Alliance DA Moderately pluralist Right Liberal
Turkey Recep Erdoğan The Justice and Development Party AKP Strongly populist Right Conservative
Selahattin Demirtaş Peoples’ Democratic Party HDP Strongly pluralist Left Liberal
Kemal Kılıçdaroğlu Republican People’s Party SHP/CHP Moderately pluralist Left Liberal
UK Nigel Farage UK Independence Party UKIP Strongly populist Right Conservative
Jeremy Corbyn Labour Party Lab Moderately populist Left Conservative
Boris Johnson Conservative Party CON Moderately populist Right Conservative
USA Donald Trump Republican Party GOP Strongly populist Right Conservative
Hilary Clinton Democratic Party Dem Moderately pluralist Left Liberal
Appendix E Supplementary: Ideological differences
We examined differences in the expression of negative emotions and neutrality of expression between different ideological groups according to the GPS variable Type_Values, which combines two categorizations of political parties: economic values (Left = pro-state or Right = pro-market) and social values (Liberal or Conservative). The results of independent samples t-test are shown in the Table 5.
Negative emotions Neutral expression
t p d t p d
Right - Left 1.447 0.151 0.203 -2.135 0.034 -0.304
Conservative - Liberal 1.715 0.088 0.214 -1.474 0.142 -0.207
Table 5: Results of the t-test for differences in mean negative emotions score and mean neutral expression between political leaders belonging to parties differing on Right VS Left and Conservative VS Liberal values according to the GPS dataset
Appendix F Supplementary: Image annotation
For the image annotation study, 514 frames were randomly selected from the sample of all frames extracted from the videos in the main study. Image annotation was performed in RStudio (Posit team 2024) using a Shiny (Chang et al. 2024) application built using a customized version of the taipan R package (Kobakian & O’Hara-Wild 2018). Figure 10 shows the main interface of the application used by coders. Responses from each of the 5 coders were exported from the application as CSV files.
Figure 10: Taipan interface showing the image from the sample (top left), short instructions (bottom left), and the questions about the image along with navigation buttons (top right).
After aggregating the results from all human coders, a majority vote principle was applied to obtain the final label. This means that at least three out of five coders had to apply the same label out of the four labels given (“No face”, “Positive”, “Negative”, “Neutral”) to the image in order for that label to be taken as the output of the human annotation process. In the case of the machine learning output, we took the highest (sum) score to obtain the final label: “Positive”, “Negative” or “Neutral”. If all scores had missing values in the case of the machine learning output, we assigned a “No face” label to that image.
Agreement between pairs of human coders and between any individual coder and machine learning output is not highly relevant to the aim of the study, however we show values of pairwise Cohen’s Kappa on Figure 11 to illustrate the difficulty of the task and the absence of high levels of agreement between human coders.
Figure 11: Pairwise Cohen’s Kappa values for agreement between pairs of human coders or pairs of human coders and the ML model.
[…]
Positive 0.5 0.25 0.23 0.02
Neutral 0.13 0.49 0.30 0.08
Negative 0.17 0.2 0.59 0.04
No Face 0.09 0.03 0.29 0.59
Precision 0.26 0.50 0.72 0.48
Recall 0.5 0.49 0.59 0.59
Table 6: Confusion matrix with row-wise ratios, precision and recall values for agreement between the majority vote of human coders and the ML model output.
In summary, when considering a 4-label problem where human labels are obtained by a majority vote, we have 458 cases with 55.2% agreement, Cohen’s Kappa of 0.31 and F1 score of 0.502. Table 6 provides additional information. The highest accuracy is recorded in the case of negative emotions and the lowest accuracy is in the case of positive emotions. This is likely due to the imbalance within emotion labels on which the ML model is trained.
Table 7 shows values for a subset of cases where the majority of humans and the ML model agree that there is a face clearly visible (present) in the image (3-label problem). Here, we have 365 cases with 59.83% agreement, Cohen’s Kappa of 0.32 and F1 score of 0.53.
[…]
Positive 0.51 0.26 0.23
Neutral 0.13 0.6 0.27
Negative 0.18 0.2 0.62
Precision 0.27 0.49 0.80
Recall 0.51 0.6 0.62
Table 7: Confusion matrix with row-wise ratios, precision and recall values for agreement between the majority vote of human coders and the ML model output in cases where there is agreement about a face being clearly visible (present)

### v5

@@ 4 Results
Results suggest that negative emotions are more frequently expressed in the entire sample, with anger (0.227) as the emotion having the highest average score, followed by sadness (0.206) and fear (0.127). As a consequence, the average proportion of negative emotions is greater than 0.5, which means that on average negative emotions are detected more frequently than positive emotions (e.g., happiness) and neutral expressions combined. There are clear issues with the detection of disgust and surprise, where a low maximum level suggests low positive predictive value for these emotions. However, our focus is on differences in the expression of negative emotions and in scores of neutral states. Aggregating negative emotions enables us to surpass the false positive issues within the group of negative emotions [e.g., expression of fear being classified as sadness, disgust as anger, 61] and minimizes the impact of the low positive predictive value of disgust and surprise.
@@
The mean average score of negative emotions in the populist group is 0.616 (95% CI 0.584–0.649), compared to the 0.500 mean average score for pluralists (95% CI 0.464–0.537). The difference between the two groups is statistically significant ($t=4.691$, $p=0.000005$, $d=0.66$ moderate effect size). This finding suggests that pluralists generally tend to have an average score of negative emotions around 0.5, which roughly translates to the situation where negative emotion is the prevailing expression in one half of the video (in fact, that percentage is 46.1% compared to 66.7% for populists; see table in the Supplement B – Dominant emotions per frame). On the other hand, for populists, more than one half of each video where they appear is dominated by negative emotions.
@@ 5 Discussion
The present study examines facial emotional expressions of populist and non-populist political leaders across fifteen countries using a deep-learning-based computer-vision algorithm to investigate differences in their expressions of negative emotions (anger, disgust, fear, and sadness) and neutral stances, i.e., facial expressions which lack emotional disposition. Our findings indicate that populist leaders tend to express negative emotions more frequently than non-populist leaders, and conversely, that neutral expressions are less common in their public performances than they are in those of more conventional, pluralist political figures. The significance and the effect size of these differences are higher compared to differences between ideological groups (see Supplement E – Ideological differences). These findings align with the theoretical framework on the characteristics of affective populist communication [5, 6, 44], as well as with previous, mostly text-based research which indicates that populist leaders employ language which is associated with negative emotion at a higher rate than mainstream, pluralist politicians [10, 62, 63, 14, 7].
@@
Overall, the results provide new evidence of a style of affective, non-verbal communication more common to political leaders whose rhetoric is populist-leaning, and thus further substantiate the idea that populism is performative in nature [20, 16]. As such, because it is subversive in relation to established political conventions, the populist performance is especially well-suited for the environment of social media where the principle of “any press is good press” often applies—that is, the quantity of audience attention bears more weight than its quality [48]. On one hand, the populist use of negative emotional displays in videos may therefore be a mechanism through which they maintain relevance, considering that controversial content online ordinarily generates more engagement among consumers [49]. On the other, populists’ facial expressions can provide affective heuristics for audiences, influencing their evaluation of candidates’ political competence [64] and allowing them to quickly reduce complex political issues into simple, emotional interpretations by visually framing them as negative.
Since this emerging dynamic has the potential to influence the behavior of voters—and perhaps even politicians themselves [^6] Due to the unmediated relationship between political figure and voter, it may very well be possible for online audiences to directly influence politicians in return. For example, once videos prominently featuring negative emotions start performing well, they might create a feedback loop which, in the long run, incentivizes political figures to “go negative” when they may have otherwise taken a different communication approach.—it holds several implications for populism, political communication, and polarization research. The Manichaen mentality of “Us versus Them" is fundamental to the populist worldview, and it further entrenches ideological positions and social divides. It is also largely dependent on fostering contentious discourse practices and negatively portraying political opponents and their beliefs. Thus, understanding how and when populist leaders bring negative emotional appeals to the forefront is vital for developing a more refined model of populist communication that would consider both the cognitive and affective sides to their political behavior. Studying non-verbal cues, in particular, can provide a breadth of information for this line of research since facial expressions of emotion add a layer of meaning to political messaging, establishing credibility and trust (or lack thereof) in political figures and appealing to audiences on a more intuitive perceptive level.
@@
Firstly, the sample of YouTube videos is not globally representative and does not provide ample coverage of the global political spectrum in regard to the use of populist or pluralistic rhetoric. As we have shown in the supplementary analysis (see Supplement C – Country-level differences), the differences between populist and pluralist leaders are not consistent in all investigated countries. Larger research infrastructure is required to process a more substantial number of videos, leaders and countries and the current computational toolset needs to be extended to enable seamless automatic processing of videos. This also relates back to the gender imbalance of the current sample. While it is reasonable to hypothesize that gender influences emotional expression among political leaders [65], in this case, including more women without significantly increasing the overall size of the sample and the number of leaders would mean focusing only on countries where there is at least one female leader of a non-fringe political party. This creates the risk of introducing bias towards Western countries in the sample and disregarding the global scope.
@@
Secondly, a key component of our analysis is the measure of populist rhetoric supplied by the Global Party Survey database [57] which significantly limits the timeframe of the analysis (no leaders, parties and videos after 2019 were taken into consideration), but also imposes the inherent limits of an expert survey. While the GPS is the only available source of multi-dimensional measurements of populist rhetoric on a global scale, the experts surveyed may operate within different understandings of what populist rhetoric is. This is particularly challenging in developing countries, where fewer experts answered the survey, as well as in autocratic regimes where experts may not have been able to provide a fair assessment of the (ruling) political parties (or refused to do so). By focusing on large countries with free elections, we tried to mitigate some of the risks, but there is the possibility that some measurements used in this study reflect expert biases, especially since we used a single, aggregate measure of populist rhetoric. Future studies should aim for more granularity and robustness in the measurement of populist rhetoric, which may be obtained by following a text-as-data approach and obtaining the measures using computational and machine learning tools, where expert survey scores provide external validation [66].
@@
Furthermore, since we lacked the resources necessary to process every political speech from each party’s official YouTube channel, the first criterion of selection is the view count of the videos. Our expectation is that this metric serves as a proxy for the relevance of the speech content to a wider audience over a longer period of time. However, since we have no information about the audience, it is possible that we are observing a demand-side effect, where audiences who consume populist videos might show a stronger preference for those that prominently feature negative emotions, introducing a specific bias in the sample.
This consideration may apply even more broadly, as our analysis does not fundamentally account for systemic bias in media consumption, coverage, and narrative construction, and consequently the differences in emotional expression may not be a result of varying political styles but rather of bias in media selection. Future research should address this issue by comparing emotional expressions across a wider range of media formats—including other social media platforms, television broadcasts of speeches and debates, mediated interviews, etc.—to determine whether the observed patterns persist beyond YouTube’s algorithm and environment.
The next major limitation is a result of the machine learning model deployed for emotion recognition. We conducted a small validation study to compare the performance of our model with the consensus of 5 human coders. Although results show evidence of moderate agreement between humans and the machine learning model (53-60%), there is high inconsistency between human coders even for the simple task of assessing whether a given facial expression shows positive emotion, negative emotion, or a neutral expression. This level of agreement, while adequate for our exploratory study, indicates substantial uncertainty in the classification of emotional expressions that could potentially affect our conclusions. The higher detection rate for negative emotions compared to positive ones could potentially amplify differences between groups if the model is more sensitive to certain types of expressions.
@@
There is an urgent need for large-scale image annotation studies with expert coders if we are to assess the performance and accuracy of FER models in the future, especially on difficult “in the wild” image datasets such as ours. However, it is worth noting that in our main analysis we focus on averages across 300 frames of each video, making our analysis robust to misclassification of a smaller number of individual frames. This methodological choice helps mitigate the impact of individual frame misclassifications on our overall findings. That being said, we believe that fair agreement between the human majority vote and machine learning model does not compromise the validity of our findings.
For future research, several technical improvements could enhance the accuracy of facial emotion recognition in political contexts. The most promising direction appears to be zero-shot classification based on open-source transformer models [67] such as OpenAI’s CLIP (Contrastive Language-Image Pre-training) [68], which may offer more nuanced performance with the ability to process facial expressions in the specific context of the image along with other visual cues. These models can leverage large-scale pre-training on diverse datasets and potentially provide more culturally sensitive interpretations of facial expressions [69]. Additionally, fine-tuning pre-trained models specifically on datasets of political figures’ expressions could improve domain-specific accuracy, as could ensemble methods that combine multiple classification approaches to enhance robustness.
@@
Overall confidence in the presented results would surely be higher if we based our analysis on a computer-vision approach which was shown to have higher accuracy, at least on FER2013 data. One of the possible alternatives would be to follow recent work [30, 70] and use proprietary API such as Amazon’s Rekognition. Nonetheless, the aim of this paper was to introduce an open-source approach focusing on free computing resources and the deployment of machine learning algorithms on systems that are accessible to individual researchers in social sciences, including graduate students without significant institutional support for computational work.
This study was focused on exploring basic mean-level differences between the given groups, but it is worth noting that the full output of our emotion detection workflow includes time series data suitable for more advanced analyses. For instance, multivariate analysis could explore affective dynamics [71], identify differences related to various factors (such as type of public address or proximity of elections), and cluster political leaders according to temporal patterns of emotional expression. Furthermore, multivariate change point detection methods [72] could identify key moments where emotional expressions significantly change. Future research could extend this approach in other promising directions as well. Network psychometric methods could model dynamic interrelationships between emotional expressions, potentially revealing different temporal patterns between populist and pluralist leaders [73]. Recent speech-to-text models like Whisper [74] could enable integrated analyses of verbal and non-verbal emotional expressions, providing insights into the alignment between spoken content and facial displays. Additionally, simulation approaches could help test theoretical predictions about emotion dynamics in political communication [75]. Collectively, these approaches would move beyond static comparisons toward analyses of temporal dynamics, potentially revealing how leaders modulate emotional displays in response to audience feedback or political contexts.
@@ 6 Data availability
Data files are stored as CSV files. Processed data used for both the main and supplementary analyses can be found at: https://github.com/atomashevic/face-of-populism. Repository contains a CSV file with a list of URLs of all YouTube videos used in the analysis as well as Python code that can be used to reproduce the CSV files with processed data. Code used for image annotation and the resulting dataset can be found at: https://github.com/atomashevic/fop-annotation.
@@ 7 Acknowledgements
The authors did not receive any financial support for this research. The annotation of images was made possible thanks to the dedicated work and effort of our graduate students, A. Nađ, M. Milanko, and A. Mišić, who volunteered as coders.